What volume of 5 M NaOH is needed to create a 100 mL solution of 1 M NaOH?
1
Expert's answer
2021-10-01T10:02:02-0400

Number of moles(n)=Molarity(M)xvolume(V) in litres

"M_1\times V_1= M_2\times V_2"

"M_1=initial \;Molarity= 5.0 \;M\\

V_1=initial\; volume = \;? \;mL\\

M_2= diluted \;Molarity = 1.00\; M\\

V_2= diluted \;volume = 100\; mL\\"


"M_1\times V_1= M_2\times V_2 \\

5.0\times V_1=1.00\times 100\\




V_1=20 mL"